# Patho-AgenticRAG: Towards Multimodal Agentic Retrieval-Augmented Generation for Pathology VLMs via Reinforcement Learning

### v1

@@ 1 Introduction
In recent years, more and more studies have attempted to introduce VLMs into digital pathology tasks, such as diagnostic assistance [1], lesion classification [2], and question-answering systems [3]. However, existing pathology VLMs still face key challenges such as severe hallucinations and a lack of structured semantic control of retrieval mechanisms, especially in tasks that require factual support and traceable evidence. Therefore, how to build a highly reliable, multimodally interpretable pathology VLM with a factual consistency assurance mechanism has become an important issue that needs to be solved in this field. Although many works in recent years have attempted to apply the Retrieval-Augmented Generation (RAG) framework to medical multimodal tasks to improve the accuracy and credibility of large language models in medical reasoning, there are still many limitations in pathological image analysis scenarios. MMed-RAG proposed a general multimodal medical RAG system [4], but the image vector library it constructed lacks fine-grained annotation of the organization system and is only divided according to the type of imaging modality (such as CT, X-ray), which cannot meet the actual needs of “classification of the organ or tissue system to which the image belongs” in pathological diagnosis. For example, in clinical work, it is necessary to clarify whether the image comes from a specific system, such as breast or lung, to match contextual knowledge and structured diagnostic paths. At the same time, some studies completely ignore the image modality and rely only on text retrieval, failing to give full play to the important role of images in assisting VLM reasoning, especially in scenarios facing image-text consistency and visual evidence support [5, 6]. In addition, recent methods such as MedRAG and Medical Graph RAG introduce complex reasoning processes guided by knowledge graphs [7, 8]. Although they enhance reasoning capabilities, the system design is complicated, the execution process lacks intelligence and scalability, and it is difficult to adapt to a variety of actual clinical tasks. Although Liu et al. emphasizes the importance of knowledge augmentation, its RAG module has the problem of insufficient instruction matching, making it difficult to stably extract relevant knowledge fragments in complex instruction following tasks, affecting model performance [9]. This study aims to build an intelligent retrieval enhancement generation framework for pathology VLMs, to improve the credibility and interpretability of the model in complex question-answering and reasoning tasks.
@@ 2.2 Reinforcement Learning for Medical VLMs
Reinforcement learning (RL) provides a promising paradigm for aligning the outputs of VLMs with clinical accuracy requirements, especially in high-risk medical domains where hallucinated descriptions can lead to severe consequences [26]. A core challenge in medical VLMs is ensuring factual alignment between visual evidence (e.g., pathology slides, radiology images) and textual outputs [27, 28].
@@ 3.1 Overall Framework
The overall architecture adopts a modular design and contains four main components. 1. Multimodal pathology knowledge base: This is a specialized vector database containing a rich collection of pathology textbook pages. It acts as an external storage for agent queries to collect relevant evidence, such as images of similar cases and their corresponding diagnostic descriptions; 2. Intelligent agentic router: This is the central processing unit of our framework. It accepts the initial diagnosis query, decomposes it into logically sequential subtasks, and plans and executes information retrieval; 3. VRAG Agent [19]: This module supports multi-turn retrieval and summarization. It interacts with the knowledge base and distills the returned textbook images into concise, useful information. 4. Core vision language model (inference engine) [26]: We use the pretrained pathology VLM as the basic inference engine. With the contextual summaries provided by the VRAG agent, it performs inference to address the diagnostic query.
@@ 3.4 Agentic Diagnostic Workflow with Multimodal Evidence Tracing
Our system adopts an intelligent multi-agent workflow that transforms a diagnostic query into an evidence-grounded conclusion. Given a user input (e.g., a pathology image and candidate diagnoses), the Agent Router module first parses the query, decomposes it into sub-tasks aligned with each diagnostic candidate, and formulates a high-level retrieval plan. It delegates retrieval and evidence aggregation to the VRAG Agent, which operates under its guidance.
@@
3.5 Tool Integrated RL for Agent Router
Traditional RAG systems are often treated as static pipelines, without adapting their behavior to each query’s complexity. To overcome this limitation, we introduce a reinforcement learning (RL) framework that enables an agent to learn dynamic invocation and routing strategies. The agent’s task is to generate a decision path specifying whether and how to call the RAG system. Formally, given an input query , the agent’s policy outputs a path , optimized to maximize the expected similarity to a ground-truth decision path
@@ 3.5 Tool Integrated RL for Agent Router
Here, is the advantage at step within each output, computed relative to other outputs for the same query. This group-wise comparison helps the agent learn from relative improvements, leading to more effective decision-making in complex reasoning paths.
@@ 4.2 Patho-AgenticRAG Evaluation Results
Closed-ended questions play a crucial role in pathology-related tasks, particularly in diagnostic classification. To evaluate model performance on such tasks, we consider two types of close-ended question datasets: (1) Yes/No questions, selected from Path-VQA and Quilt-VQA; and (2) multiple-choice questions, sourced from PathMMU [36], MedXpertQA [37], and OmniMedVQA [38].
The results on close-ended benchmarks are summarized in tables 2 and 3. Patho-AgenticRAG achieves the best overall performance across most tasks, significantly outperforming both general-purpose vision-language models (e.g., InternVL3, Qwen2.5VL) and domain-specialized baselines such as Patho-R1-7B [26]. Specifically, Patho-AgenticRAG achieves +13.37% improvement on Quilt-VQA (75.80% vs. 64.72%) and +38.00% on MedXpertQA (60.00% vs. 22.00%) over Patho-R1. The largest margin appears on MedXpertQA, highlighting the importance of retrieval-augmented reasoning in knowledge-intensive tasks. On OmniMedVQA Bright Challenge, the model improves from 70.79% (Patho-R1) to 90.11%, a +19.32% increase, demonstrating substantial gains in both generalization and diagnostic precision.

### v2

@@ 1 Introduction
In recent years, more and more studies have attempted to introduce VLMs into digital pathology tasks, such as diagnostic assistance [1], risk stratification [2], and question-answering systems [3]. However, existing pathology VLMs still face key challenges such as severe hallucinations and a lack of structured semantic control of retrieval mechanisms, especially in tasks that require factual support and traceable evidence. Therefore, how to build a highly reliable, multimodally interpretable pathology VLM with a factual consistency assurance mechanism has become an important issue that needs to be solved in this field. Although many works in recent years have attempted to apply the Retrieval-Augmented Generation (RAG) framework to medical multimodal tasks to improve the accuracy and credibility of large language models in medical reasoning, there are still many limitations in pathological image analysis scenarios. MMed-RAG proposed a general multimodal medical RAG system [4], but the image vector library it constructed lacks fine-grained annotation of the organization system and is only divided according to the type of imaging modality (such as CT, X-ray), which cannot meet the actual needs of “classification of the organ or tissue system to which the image belongs” in pathological diagnosis. For example, in clinical work, it is necessary to clarify whether the image comes from a specific system, such as breast or lung, to match contextual knowledge and structured diagnostic paths. At the same time, some studies completely ignore the image modality and rely only on text retrieval, failing to give full play to the important role of images in assisting VLM reasoning, especially in scenarios facing image-text consistency and visual evidence support [5, 6]. In addition, recent methods such as MedRAG and Medical Graph RAG introduce complex reasoning processes guided by knowledge graphs [7, 8]. Although they enhance reasoning capabilities, the system design is complicated, the execution process lacks intelligence and scalability, and it is difficult to adapt to a variety of actual clinical tasks. Although Liu et al. emphasizes the importance of knowledge augmentation, its RAG module has the problem of insufficient instruction following, making it difficult to stably extract relevant knowledge fragments in complex instruction following tasks, affecting model performance [9]. This study aims to build an intelligent retrieval augmented generation framework for pathology VLMs, to improve the credibility and interpretability of the model in complex question-answering and reasoning tasks.
@@ 2.2 Reinforcement Learning for Medical VLMs
Reinforcement learning (RL) provides a promising paradigm for aligning the outputs of VLMs with clinical accuracy requirements, especially in high-risk medical domains where hallucinated descriptions can lead to severe consequences [26]. A core challenge in medical VLMs is ensuring factual alignment between visual evidence (e.g., pathology slides, radiology images) and textual outputs [27, 28, 29].
@@ 3.1 Overall Framework
The overall architecture adopts a modular design and contains four main components. 1. Multimodal pathology knowledge base: This is a specialized vector database containing a rich collection of pathology textbook pages. It acts as an external storage for agent queries to collect relevant evidence, such as images of similar cases and their corresponding diagnostic descriptions; 2. Intelligent agentic router: This is the central processing unit of our framework. It accepts the initial diagnosis query, decomposes it into logically sequential subtasks, and plans; 3. VRAG Agent [19]: This module supports multi-turn retrieval and summarization. It interacts with the knowledge base and distills the returned textbook images into concise, useful information. 4. Core vision language model (inference engine) [26]: We use the pretrained pathology VLM as the basic inference engine. With the contextual summaries provided by the VRAG agent, it performs inference to address the diagnostic query.
@@ 3.4 Agentic Diagnostic Workflow with Multimodal Evidence Tracing
Our system adopts an intelligent multi-agent workflow that transforms a diagnostic query into an evidence-grounded conclusion. Given a user input (e.g., a pathology image and candidate diagnoses), the Agentic Router module first parses the query, decomposes it into sub-tasks aligned with each diagnostic candidate, and formulates a high-level retrieval plan. It delegates retrieval and evidence aggregation to the VRAG Agent, which operates under its guidance.
@@
3.5 Tool Integrated RL for Agentic Router
Traditional RAG systems are often treated as static pipelines, without adapting their behavior to each query’s complexity. To overcome this limitation, we introduce a reinforcement learning (RL) framework that enables an agent to learn dynamic invocation and routing strategies [35]. The agent’s task is to generate a decision path specifying whether and how to call the RAG system. Formally, given an input query , the agent’s policy outputs a path , optimized to maximize the expected similarity to a ground-truth decision path
@@ 3.5 Tool Integrated RL for Agentic Router
Here, is the advantage at step within each output, computed relative to other outputs for the same query. This group-wise comparison helps the agent learn from relative improvements, leading to more effective decision-making in complex reasoning paths. See Appendix C for details.
@@ 4.2 Patho-AgenticRAG Evaluation Results
Closed-ended questions play a crucial role in pathology-related tasks, particularly in diagnostic classification. To evaluate model performance on such tasks, we consider two types of close-ended question datasets: (1) Yes/No questions, selected from Path-VQA [38] and Quilt-VQA [39]; and (2) multiple-choice questions, sourced from PathMMU [40], MedXpertQA [41], and OmniMedVQA [42].
The results on close-ended benchmarks are summarized in Tables 2 and 3. Patho-AgenticRAG achieves the best overall performance across most tasks, significantly outperforming both general-purpose vision-language models (e.g., InternVL3, Qwen2.5VL) and domain-specialized baselines such as Patho-R1-7B [26]. Specifically, Patho-AgenticRAG achieves +13.37% improvement on Quilt-VQA (75.80% vs. 64.72%) and +38.00% on MedXpertQA (60.00% vs. 22.00%) over Patho-R1. The largest margin appears on MedXpertQA, highlighting the importance of retrieval-augmented reasoning in knowledge-intensive tasks. On OmniMedVQA Bright Challenge, the model improves from 70.79% (Patho-R1) to 90.11%, a +19.32% increase, demonstrating substantial gains in both generalization and diagnostic precision. Details are in Appendix D
@@ 5 Conclusion
Appendix A Multimodal Knowledge Base Construction
To construct a high-quality knowledge base tailored for pathology, we collected and processed over 600 authoritative pathology textbooks that span a wide range of diagnostic domains. These sources were chosen for their depth, credibility, and relevance to real-world diagnostic practice.
To improve retrieval accuracy and reduce noise in downstream reasoning, we removed sections that typically lack diagnostic content, including covers, prefaces, tables of contents, and references. The remaining pages, primarily rich in textual explanations and pathology figures, were preserved as the core knowledge units. These preserved pages were further categorized into 19 distinct anatomical and diagnostic classes to support systematized retrieval. Specifically, the categories are as follows:
1.
Bone and soft tissue – including neoplastic and non-neoplastic lesions of bone, cartilage, and connective tissue.
2.
Cytology – encompassing exfoliative and fine-needle aspiration cytopathology across body systems.
3.
Gastrointestinal tract, liver, gallbladder, pancreas, and digestive system – covering both luminal and solid organ pathology within the digestive system.
4.
Hematology, lymphatic system, and bone marrow – focusing on hematopoietic and lymphoid neoplasms and marrow disorders.
5.
Infectious diseases – including pathology of viral, bacterial, fungal, and parasitic infections.
6.
Oral and head and neck – covering salivary glands, oral mucosal lesions, larynx, pharynx, nasal cavity, and other ENT (ear, nose, throat) structures.
7.
Urinary and male reproductive system – involving kidney, bladder, prostate, testis, and related structures.
8.
Breast – dedicated to benign, precursor, and malignant breast lesions.
9.
Endocrine system – covering thyroid, parathyroid, adrenal glands, and neuroendocrine tumors.
10.
General comprehensive pathology – consisting of cross-system concepts such as inflammation, necrosis, and neoplasia.
11.
Histology and embryology – including normal tissue architecture and developmental processes.
12.
Neonatal, pediatric, and childhood diseases – covering age-specific pathologies from birth to adolescence.
13.
Skin and dermatopathology – addressing inflammatory, infectious, and neoplastic skin conditions.
14.
Central nervous system – including brain, spinal cord, and peripheral nerve pathology.
15.
Female reproductive system – encompassing ovary, uterus, cervix, and vulvovaginal tract pathology.
16.
Gross specimen sampling – providing guidance on macroscopic examination and tissue sectioning.
17.
Immunohistochemistry and molecular pathology – focused on biomarker detection and molecular diagnostics.
18.
Ophthalmology and otolaryngology – covering diseases of the eye, ear, and related sensory organs.
19.
Trachea, lung, pleura, respiratory system, and mediastinum – including neoplastic and inflammatory diseases of the thoracic cavity.
Each cleaned PDF was converted page-wise into an image format, ensuring that the visual structure (including images, captions, and layout) remained intact.
We used the ColQwen2 model [32], a powerful vision-language encoder, to generate unified high-dimensional embeddings for each image-text pair. This model jointly encodes both the fine-grained visual features and the contextual diagnostic semantics into a shared vector space, enabling effective multimodal similarity retrieval.
All embeddings were stored in Milvus [34], an open-source vector database optimized for large-scale similarity search. To accelerate retrieval in high-dimensional space, we employed the hierarchical navigable small world (HNSW) indexing algorithm [33], which is well suited for fast, approximate nearest neighbor search at scale. This infrastructure allows our reasoning agents to retrieve highly relevant visual evidence from hundreds of thousands of multimodal entries in real time.
Appendix B Details of the Agentic Diagnostic Workflow
The intelligent diagnostic pipeline operates in four distinct stages, involving three key modules: the Agentic Router (task planner), the VRAG Agent (retrieval and summarization), and the Pathology VLM (inference engine). Each step is described in detail below.
Step 1: Query Understanding and Task Decomposition (Agentic Router) The Agentic Router receives the user’s input and identifies the semantic structure of the task, typically involving image understanding and candidate differentiation. It decomposes the query into subproblems mapped to each candidate diagnosis.
Step 2: Entity-Centric Multimodal Retrieval (within VRAG Agent) For each candidate diagnosis, the VRAG Agent performs a two-step retrieval process:
Textual Retrieval: Using candidate-specific keywords, the system retrieves the top 20 textual entries from the knowledge base.
Multimodal Re-ranking: Retrieved entries are then ranked by calculating joint similarity between the input image and associated image-text pairs using the ColQwen2 embedding [32]. The similarity score is computed according to Equation 1, which captures the cross-modal relevance between the query image and each retrieved image-text pair in the embedding space. The highest-ranked entry becomes the evidence for that candidate.
Step 3: Iterative Evidence Aggregation (VRAG Agent) The VRAG Agent [19] performs multi-turn retrieval and refinement. It dynamically evaluates evidence sufficiency, performs focused re-retrieval when necessary, and aggregates the most discriminative content into a structured prompt. The prompt includes: (1) the original user input, (2) top retrieved evidence for each diagnosis, and (3) explicit comparative instructions.
Step 4: Reasoning and Report Generation (Pathology VLM) The structured prompt is passed to a specialized vision-language model fine-tuned for diagnostic tasks [26]. The model conducts contrastive reasoning, constrained to the provided evidence, and returns both a predicted diagnosis and an interpretable justification.
Appendix C Details of Models and Training
C.1 Training Dataset Construction
The training data for the agentic router was derived from the training sets of Quilt-VQA[39] and Path-VQA[38]. We first ran inference using Patho-R1 and filtered the samples based on the correctness of its responses. Specifically, we removed questions with overly simple reasoning or insufficient thought chains, as they were not useful for training our agent flow. From the remaining samples, we randomly selected 2,200 questions that Patho-R1 answered incorrectly (indicating a need for RAG) and 2,200 that it answered correctly (indicating no need for RAG) to construct the training dataset. Among these selected questions, we ensured an even split between MCQ and YORN formats. We then distilled the ground truth using QwenMax, ultimately retaining 4,000 samples for RL training and an additional 400 samples for SFT training, as shown in Figure 5.
Specifically, to construct ground-truth decision paths for reinforcement learning, we prompted QwenMax with carefully designed instructions that emulate expert diagnostic reasoning. Specifically, given a user query , we provided QwenMax with a prompt that specifies: (1) available tools (e.g., rag, classifier), (2) a comprehensive list of anatomical partitions, and (3) detailed behavioral guidelines for tool invocation.
The model was instructed to first output a <think> field capturing its reasoning, followed by one or more structured <tool_call> entries. For instance, it would decide whether to query external knowledge via the rag tool, invoke the classifier tool to infer tissue system based on question content, or rely solely on internal knowledge when appropriate. A sample prompt template and tool call schema are provided in Appendix D.
The resulting outputs serve as high-quality ground-truth action trajectories, reflecting both whether to invoke the retrieval system and how to structure the tool usage. These trajectories form the supervision signal for training the agent’s policy , which aims to recover the expert-like decision plan from an input query.
C.2 Training Details
All tool calls in the dataset are represented in JSON format. We use Qwen3-4B as the base model for all experiments, and training was conducted on 8 NVIDIA RTX 4090 GPUs.
Supervised fine-tuning: We adopted the LLaMA-Factory111https://github.com/hiyouga/LLaMA-Factory.git framework and froze the vision tower. We used a learning rate of 1e-5 and trained on 400 samples for 3 epochs.
Reinforcement learning: We adopted the verl222https://github.com/volcengine/verl framework for reinforcement learning. We set the actor and critic learning rates 1e-6 and 1e-5 respectively. For GRPO, we trained on 4k samples for 3 epochs
Appendix D Experimental Benchmarks, Evaluation Metrics, and Prompts
We presented the prompts used throughout our Patho-AgenticRAG workflow, along with the benchmarks and evaluation protocols adopted for assessment.
D.1 Multimodal Benchmarks for Pathology
We evaluate our approach using multiple multimodal benchmarks relevant to pathology. For the multiple-choice (MCQ) setting, we adopt the PathMMU dataset, a pathology-specific benchmark comprising images and expert-annotated questions. Following the dataset’s protocol, we downloaded pathology images originally shared via Twitter. However, due to post deletions, some images were no longer accessible, and the corresponding questions were removed from evaluation.
The PathMMU-test-tiny split includes a total of 1,139 questions, distributed as follows: Atlas (208), EduContent (255), PathCLS (177), PubMed (281), and SocialPath (218). The full PathMMU-test split contains 8,454 questions: Atlas (799), EduContent (1,683), PathCLS (1,632), PubMed (2,787), and SocialPath (1,553).
To broaden evaluation coverage, we further curated pathology-focused subsets from general medical VQA datasets. Specifically, we selected 90 pathology-related examples from MedXpertQA, and used the BRIGHT Challenge subset (890 cases) from OmniMedVQA, which focuses on diagnostic reasoning across medical specialties.
Finally, for the YORN dataset, we collected closed-ended questions from the test splits of Path-VQA and Quilt-VQA, resulting in 3,362 and 343 questions, respectively.
D.2 Evaluation Metrics for Multi-Model Fusion
We evaluate retrieval performance using three metrics: mean reciprocal rank at (MRR@), recall@, and normalized discounted cumulative gain at (NDCG@), defined as follows:
Recall@ measures the fraction of queries for which the target textbook page is included among the top retrieved results. Recall@ is given by
(9) where is the indicator function.
Mean Reciprocal Rank at (MRR@) reports the average reciprocal rank of the target textbook page within the top results. Let be the number of queries and the position of the target textbook page for query . The metric is defined as
(10) where the reciprocal rank is zero if the target textbook page does not appear in the top .
Normalized Discounted Cumulative Gain at (NDCG@) evaluates ranking quality based on the graded relevance of retrieved textbook pages. For NDCG calculation, the target textbook page is assigned a relevance score of , and the previous and next textbook pages are assigned a score of , reflecting that neighboring pages may contain partially relevant information. NDCG@ is computed as
(11) where
(12) and is the relevance score of the -th retrieved textbook page for query . The ideal DCG, , is computed with pages sorted by the highest possible relevance.
D.3 Evaluation Metrics for Medical QA Tasks
To evaluate model performance on various pathology benchmarks, we adopt standard accuracy-based metrics. For closed-ended visual question answering tasks, including multiple-choice and yes-or-no questions, we compute exact-match accuracy, where a prediction is considered correct only if it exactly matches the ground-truth answer. In addition, we report per-category scores on datasets such as PathMMU, whose subsets originate from diverse real-world sources (e.g., textbooks, educational slides, social media, and literature). This allows us to assess the model’s generalization ability across heterogeneous pathology domains.